Solve each quadratic in form equation.

9. x^4-8x^2+7=0

10. x^4-2x^2-35=0
1
Expert's answer
2013-01-17T09:29:13-0500
9.& x^4-8x^2+7=0
x^2=y
y^2-8y+7=0
D=64-4*7*1=36
y=(8+6)/2= 7
y=(8-6)/2=1
x1= sqrt(7)
x2=-sqrt(7)
x3=1
x4=-1

10. x^4-2x^2-35=0
x^2=y
y^2-2y-35=0
D=4+4*35*1=144
y=(2+12)/2= 7
y=(2-12)/2=irrational roots
x1= sqrt(7)
x2=-sqrt(7)